OpenStudy (anonymous):

Write four different algebraic expressions to be translated. (ie 2x+4) Examples must include variables, and there must be at least one example with an addition, subtraction, multiplication, division, parenthesis and exponent, and one with a combination of phrases or symbols. 1. 2. 3. 4.

OpenStudy (driftracer305):

y = 2x^2 + 3x -4 y = 2(3-4)^2 + 5 y = 2/7 x +12 -2x-y = 1

OpenStudy (driftracer305):

ohh wait........i get it One-sixth of f is added to 6 = 1/6f + 6 Add two-thirds to 9 times s = 9s + 2/3 One-fourth of y is added to the product of 6 and g = 1/4y + 6g Subtract 5 from 6 times k = 6k-5

OpenStudy (driftracer305):

u can use both buddy........ :D